### v1

# Weak* decomposition and Radon-Nikodym theorem for quantum expectations

Abstract
A quantum expectation is a positive linear functional of norm one on a non-commutative probability space (i.e., a C*-algebra). For a given pair of quantum expectations $\mu$ and $\lambda$ on a non-commutative probability space $A$, we propose a definition for weak* continuity and weak* singularity of $\mu$ with respect to $\lambda$. Then, using the theory of von Neumann algebras, we obtain the natural weak* continuous and weak* singular parts of $\mu$ with respect to $\lambda$. If $\lambda$ satisfies a weak tracial property known as the KMS condition, we show that our weak* decomposition coincides with the Arveson-Gheondea-Kavruk Lebesgue (AGKL) decomposition. This equivalence allows us to compute the Radon-Nikodym derivative of $\mu$ with respect to $\lambda$. We also discuss the possibility of extending our results to the positive linear functionals defined on the Cuntz-Toeplitz operator system.
1 Introduction
By the Riesz-Markov theorem and the Gelfand representation of unital commutative C*-algebras, one might refer to any positive linear functional (PLF) on a general C*-algebra as a non-commutative measure. However, this interpretation is not exact. In fact, the actual counterparts of classical measures are weights on von Neumann algebras, which are only densely defined and maybe unbounded. On the other hand, by Voiculescu’s non-commutative probability [7], every member of a C*-algebra is a random variable, and applying a unital PLF against such a random variable should reveal its expected value. So, for a PLF on a C*-algebra, the term quantum (or non-commutative) expectation is more reasonable than a non-commutative measure. In addition, for any PLF $\mu$ on a unital C*-algebra $A$, we have $\|\mu\|=\mu(1)$, so we can rescale $\mu$ to be a quantum expectation on $A$. Thus, studying PLFs on C*-algebras is equivalent to studying quantum expectations on them. Despite this subtle distinction in terminology, we can still adapt and reconstruct certain concepts from classical measure theory within the non-commutative framework. We should mention that PLFs on C*-algebras are important for quantum operations and quantum information theory, see [9].
Here, we consider a unital C*-algebra $A$ equipped with a positive linear functional (PLF) $\lambda$. Next, we try to compare any other PLF $\mu$ on $A$ with respect to $\lambda$. In particular, we propose a definition for weak* continuity and weak* singularity of $\mu$ with respect to $\lambda$. Then, by the theory of von Neumann algebras, we obtain a natural weak* continuous and weak* singular decomposition of $\mu$ with respect to $\lambda$. If $\lambda$ has some kind of weak tracial property known as the KMS condition, we show that our weak* method is equivalent to Arveson-Gheondea-Kavruk Lebesgue (AGKL) decomposition. The equivalence with AGKL decomposition helps us to calculate the Radon Nikodym derivative of $\mu$ with respect to $\lambda$. Clearly, one can use our approach to compare quantum expectations on a non-commutative probability space and obtain the relative Radon-Nikodym derivative of them.
2 Preliminaries
We recall some fundamental results from Banach algebra theory that will be used in proofs without explicit mention.
Definition 1.
Let $A$ be a Banach algebra, and $X$ be a Banach space. We call $X$ a left Banach $A$-Module if $X$ is an algebraic module over $A$, and if the bilinear map
$A\times X$$\longrightarrow$ $X$
$(a,x)$ $\longrightarrow a.x$
is bounded, i.e., $||a.x||\leq C||a||||x||$. Right Banach $A$-modules are defined similarly. $X$ is called a Banach $A$-bimodule, if it is both left and right $A$-module plus
$a.(x.b)=(a.x).b,\quad a,~{}b\in A,~{}x\in X;$
that is, we have certain associativity between left and right actions. The module $X$ is unit linked if $A$ is unital and $1.x=x$ for any $x\in X$.
Remark 1.
When $X$ is a left $A$-module, we can make $X^{*}$ into a right $A$-module via
$X^{*}\times A$$\longrightarrow$ $X^{*}$
$(f,a)$$\longrightarrow$ $f.a~{},\qquad\langle f.a,~{}x\rangle=\langle f,~{}a.x\rangle.$
In this case, we call $X^{*}$ a dual Banach $A$-module. In addition, we can make $X^{**}$ into a left $A$-module again by declaring
$A\times X^{**}$$\longrightarrow$ $X^{**}$
$(a,F)$$\longrightarrow$ $a.F~{},\qquad\langle a.F,~{}f\rangle=\langle F,~{}f.a\rangle.$
Similarly, we can start with a right $A$-module $X$ and proceed as above, and define left and right actions on $X^{*}$ and $X^{**}$ respectively. $\blacksquare$
One instance of the above situation is when $A$ acts on itself, i.e., $X=A$, through its multiplication, which leads to the Arens’ definition of product on the second dual of any Banach algebra.
Definition 2.
Let $A$ be Banach algebra and $A^{**}$ be its second dual. For any $\Phi,~{}\Psi\in A^{**}$, we denote the first Arens product of $\Phi$ and $\Psi$ by $\Phi\square\Psi$ and define it by :
$\langle\Phi\square\Psi,~{}f\rangle=\langle\Phi,~{}\Psi\cdot f\rangle,~{}f\in A^{*}$
$\langle\Psi\cdot f,~{}a\rangle=\langle\Psi,~{}f\cdot a\rangle,~{}a\in A$
$\langle f\cdot a,~{}b\rangle=\langle f,~{}ab\rangle,~{}b\in A.$
We also we denote the second Arens product of $\Phi$ and $\Psi$ by $\Phi\diamond\Psi$ and define it by :
$\langle\Phi\diamond\Psi,~{}f\rangle=\langle\Psi,~{}f\cdot\Phi\rangle,~{}f\in A^{*}$
$\langle f\cdot\Phi,~{}a\rangle=\langle\Phi,~{}a\cdot f\rangle,~{}a\in A$
$\langle a\cdot f,~{}b\rangle=\langle f,~{}ba\rangle,~{}b\in A.$
Remark 2.
  1. 1.
    Consider the natural embedding $j:A\longrightarrow A^{**}$ and put the first Arens product on $A^{**}$. For any $a,~{}b\in A$, let $\Phi=j(a)$ and $\Psi=j(b)$. By the first Arens product, for any $f\in A^{*}$ we have that
    $\langle j(a)\square j(b),f\rangle=\langle f,ab\rangle=\langle j(ab),f\rangle;$
    hence, $j:A\longrightarrow A^{**}$ is a homomorphism if we endow $A^{**}$ with its first Arens product.
  2. 2.
    Alternatively, we can define the first and second Arens product by the aid of the natural embedding $j:A\longrightarrow A^{**}$. Note that by the Goldstine theorem $\overline{j(A)}^{wk*}=A^{**}$, where the weak* topology is $\sigma(A^{**},A^{*})$ topology. Now, if we let $\Phi=\mathrm{w^{*}lim}_{\alpha}j(a_{\alpha})$ and $\Psi=\mathrm{w^{*}lim}_{\beta}j(b_{\beta})$, then
    $\langle\Phi\square\Psi,~{}f\rangle=\mathrm{w^{*}lim}_{\alpha}\mathrm{w^{*}lim}_{\beta}\langle j(a_{\alpha}b_{\beta}),~{}f\rangle,~{}~{}f\in A^{*}$
    $\langle\Phi\diamond\Psi,~{}f\rangle=\mathrm{w^{*}lim}_{\beta}\mathrm{w^{*}lim}_{\alpha}\langle j(a_{\alpha}b_{\beta}),~{}f\rangle,~{}~{}f\in A^{*}$
  3. 3.
    Now comes the question whether the Arens first and second product on the second dual $A^{**}$ of a Banach algebra $A$ are equal. If they coincide, we call $A$ Arens regular; otherwise, we call it Arens irregular. One can show that for a locally compact group $G$, the convolution Banach algebra $L^{1}(G)$ is Arens regular if and only if $G$ is finite [19, Example 5.1.19]. This is while that any operator algebra is Arens regular [1, Corollar 2.5.4].
  4. 4.
    Finally, we should mention that if a Banach algebra $A$ is Arens regular, then the Arens product on $A^{**}$ is separately weak* continuous [1, Remark 2.5.3]. $\blacksquare$
The following definition is taken from [7] and [22].
Definition 3.
A non-commutative probability space is a unital operator algebra $A$ equipped with a unital bounded linear functional $\phi$. The functional $\phi$ is referred to as the non-commutative (or quantum) expectation, and the elements of $A$ are called non-commutative (or quantum) random variables. If $A$ is a C*-algebra or a von Neumann algebra, we additionally require $\phi$ to be a state or a normal state, respectively.
Remark 3.
Let $A$ be a unital C*-algebra and $B$ be a C*-sub-algebra of $A$ containing the unit element of $A$. By [15, Corollary 3.3.4], any positive linear functional on any unital C*-algebra attains its norm at the identity element. Thus, we have the following observations:
  1. (i)
    Suppose that $\mu:A\longrightarrow\mathbb{C}$ is a positive linear functional on $A$, and $\mu^{\prime}:B\longrightarrow\mathbb{C}$ is the restriction of $\mu$ to $B$. Then, $\|\mu\|=\|\mu^{\prime}\|$.
  2. (ii)
    Conversely, suppose that $\nu:B\longrightarrow\mathbb{C}$ is a positive linear functional on $B$, and $\tilde{\nu}:A\longrightarrow\mathbb{C}$ is a positive linear functional extending $\nu$ to $A$. Then, $\|\tilde{\nu}\|=\|\nu\|$.
Definition 4.
A C*-dynamical system $(\mathscr{A},G,\alpha)$ consists of a C*-algebra $\mathscr{A}$, a locally compact group $G$, and a $\ast$-homomorphism $\alpha:G\to\operatorname{Aut}(\mathscr{A})$ such that the action
$G\times\mathscr{A}\longrightarrow\mathscr{A},\quad(t,a)\mapsto\alpha_{t}(a),$
is continuous with respect to the usual topologies of $G$ and $\mathscr{A}$. A covariant representation of $(\mathscr{A},G,\alpha)$ is a pair $(\pi,\rho)$, where
  1. 1.
    $\pi$ is a $\ast$-representation of $\mathscr{A}$ into $\mathscr{B}\left(\mathscr{H}\right)$
  2. 2.
    $\rho$ is a unitary representation of $G$ into $\mathscr{B}\left(\mathscr{H}\right)$
  3. 3.
    The “covariant covariant ( or intertwining ) relation” holds, i.e.,
    $\pi(\alpha_{t}(a))=\lambda(t)\pi(a)\lambda(t)^{*}.$
Definition 5.
A W*-dynamical system $(\mathscr{A},G,\alpha)$ consists of a von Neumann algebra $\mathscr{A}$ contained in some $\mathscr{B}(\mathscr{H})$, a locally compact group $G$, and a $\ast$-homomorphism $\alpha$ from $G$ into the $\ast$-automorphisms of $\mathscr{A}$ such that the action $G\times\mathscr{A}\longrightarrow\mathscr{A}$ given by $(t,a)\mapsto\alpha_{t}(a)$ is continuous when $G$ is equipped with its topology and $\mathscr{A}$ is endowed with the strong operator topology (SOT). Recall that we have the natural identification
$L^{2}(G)\otimes_{2}\mathscr{H}\cong L^{2}(G,\mathscr{H}).$
A covariant representation of the W*-dynamical system $(\mathscr{A},G,\alpha)$ is a pair $(\pi,\lambda)$, where:
  1. 1.
    $\pi$ is a $\ast$-representation of $\mathscr{A}$ into $\mathscr{B}\left(L^{2}(G)\otimes_{2}\mathscr{H}\right)$ given by
    $(\pi(a)f)(s)=\alpha_{s}(a)f(s),$
    for all $a\in\mathscr{A}$, $f\in L^{2}(G,\mathscr{H})$, and $s\in G$.
  2. 2.
    $\lambda$ is a unitary representation of $G$ into $\mathscr{B}\left(L^{2}(G)\otimes_{2}\mathscr{H}\right)$ defined by
    $(\lambda(t)f)(s)=f(t^{-1}s).$
  3. 3.
    The covariant ( or intertwining ) relation holds, i.e.,
    $\pi(\alpha_{t}(a))=\lambda(t)\pi(a)\lambda(t)^{*}.$
The W*-crossed product $\mathscr{A}\overline{\rtimes}_{\alpha}G$ is defined as the von Neumann algebra generated by the images of $\pi(\mathscr{A})$ and $\lambda(G)$ inside $\mathscr{B}\left(L^{2}(G)\otimes_{2}\mathscr{H}\right)$, that is,
$\mathscr{A}\overline{\rtimes}_{\alpha}G=\left[\pi(\mathscr{A})\cup\lambda(G)\right]^{{}^{\prime\prime}}.$
3 The weak* decomposition of quantum expectations
The following observation is crucial for constructing our decomposition, so we state it as a remark for later use.
Remark 4.
Suppose that $\lambda$ is a classical , finite, positive, Radon measures on the unit circle $\mathbb{T}$. Consider the Laurent transformation
$\pi:L^{\infty}(\mathbb{T},\lambda)$$\longrightarrow$ $\mathscr{B}(L^{2}(\mathbb{T},\lambda))$
$\pi(f)$ $=$ $M_{f},\quad M_{f}(g)=fg,$ (1)
which is an isometric unital $\ast$-isomorphism from $L^{\infty}(\mathbb{T},\lambda)$ onto the range of $\pi$, see [15, Example 2.5.1]. By [6, Proposition 4.51], the weak* topology of $L^{\infty}(\mathbb{T},\lambda)$ coincides with the weak operator topology of $\pi(L^{\infty}(\mathbb{T},\lambda))$. Since $\lambda$ is a Radon measure, $C(\mathbb{T})$ can be regarded as a closed subspace (even sub-C*-algebra) of $L^{\infty}(\mathbb{T},\lambda)$, see [8, P. 184 and Section 7.2]. By [6, Corollary 4.53], von-Neumann double commutant theorem [15, Theorem 4.1.5], and [15, Corollary 4.2.8], we have
$\overline{\pi(C(\mathbb{T}))}^{wk*}=\pi(C(\mathbb{T}))^{\prime\prime}=L^{\infty}(\mathbb{T},\lambda)$
Now, let $\mu$ be an arbitrary classical, finite, positive, Radon measure on the unit circle. By Riesz-Markov Theorem [8, Corollary 7.18] and [20, Theorem III.1.1], $\mu$ can be regarded as a positive linear functional (PLF) on $C(\mathbb{T})$ given by integration against $\mu$, i.e.,
$\mu:C(\mathbb{T})$$\longrightarrow \displaystyle\mathbb{C}$
[…]
By [20, Theorem III.1.2], we see that the GNS construction of $\lambda$ is given by the mapping 4, i.e.,
$\pi_{\lambda}:C(\mathbb{T})$$\longrightarrow$ $\mathscr{B}(L^{2}(\mathbb{T},\lambda))$
$\pi_{\lambda}(f)$ $=$ $M_{f},\quad M_{f}(g)=fg.$ (2)
So, in fact
$\overline{\pi_{\lambda}(C(\mathbb{T}))}^{wk*}=\pi_{\lambda}(C(\mathbb{T}))^{\prime\prime}=L^{\infty}(\mathbb{T},\lambda)$ (3)
On the other hand, $L^{\infty}(\mathbb{T},\lambda)$ is a von Neumann algebra with the pre-dual $L^{1}(\mathbb{T},\lambda)$. By the general theory of von Neumann algebras [20, Sections II.2 and III.2], $L^{1}(\mathbb{T},\lambda)$ is the space of all weak* continuous linear functionals on $L^{\infty}(\mathbb{T},\lambda)$. With this preamble in mind, we would like to explore the absolute continuity of a measure $\mu$ with respect to a measure $\lambda$ and its relationship with weak* continuous functionals on $L^{\infty}(\mathbb{T},\lambda)$. In particular, we will show that for a positive function $f\in L^{1}(\mathbb{T},\lambda)$, the measure or PLF $fd\lambda$ defined by
$fd\lambda:C(\mathbb{T})$$\longrightarrow \displaystyle\mathbb{C}$
[…]
can be extended to the PLF
$fd\lambda:L^{\infty}(\mathbb{T},\lambda)$$\longrightarrow \displaystyle\mathbb{C}$
[…]
which defines a weak* continuous on $L^{\infty}(\mathbb{T},\lambda)$. Interestingly, these are all the possible weak* functionals, and by the Radon-Nikodym theorem, they are all absolutely continuous with respect to $\lambda$. $\blacksquare$
The following classical observation seems to be a well-known theorem and is crucial for our weak* decomposition of PLFs. However, we could not find a specific reference for it. Thus, we state it explicitly and provide a proof based on the preceding remark
Theorem 1.
Let $\mu$ and $\lambda$ be two positive classical Radon measures on the unit circle, and by Remark 4 consider them as positive linear functionals on $C(\mathbb{T})$. Then,
  1. (i)
    $\mu$ can be extended to a weak*-continuous functional on ${L}^{\infty}(\lambda)$ if and only if $\mu$ is absolutely continuous with respect to $\lambda$.
  2. (ii)
    No extension of $\mu$ to ${L}^{\infty}(\lambda)$ can majorize a non-zero weak*-continuous PLF on ${L}^{\infty}(\lambda)$ if and only if $\mu$ is singular with respect to $\lambda$.
  3. (iii)
    $\mu$ can be decomposed with respect to $\lambda$ into two parts, a weak* continuous part and a weak* singular part, i.e., $\mu=\mu_{wc}+\mu_{ws}.$
Proof.
  1. (i)
    Suppose that $\mu$ and $\lambda$ are two classical positive Radon measures on $\mathbb{T}$. By Riesz-Markov theorem, $\mu:C(\mathbb{T})\longrightarrow\mathbb{C}$ is a PLF acting by integration against $\mu$. Let $\mu$ has a weak-continuous extension $\hat{\mu}$ to a PLF on $L^{\infty}(\mathbb{T},\lambda)=\overline{C(\mathbb{T})}^{wk*}$. Here, wk* is the the weak* topology of $L^{\infty}(\mathbb{T},\lambda)=(L^{1}(\mathbb{T},\lambda))^{*}$. By the definition of weak* topology, any weak* continuous functional corresponds to a member of $L^{1}(\mathbb{T},\lambda)$. By Remark 4, we can consider $L^{\infty}(\mathbb{T},\lambda)$ inside $\mathscr{B}(L^{2}(\lambda))$ as multiplication by bounded operators such that weak* topology of $L^{\infty}(\mathbb{T},\lambda)$ coincides with the weak operator topology it receives from $\mathscr{B}(L^{2}(\lambda))$. Since the measures in question are positive, and $\hat{\mu}$ is continuous with respect to the weak operator topology, by [15, Theorem 4.2.6], we can find a positive $f\in L^{1}(\mathbb{T},\lambda)$ such that for any $g\in C(\mathbb{T})$
[…]
    Hence, $\mu$ and $fd\lambda$ give rise to the same functional on $C(\mathbb{T})$, i.e., $d\mu=fd\lambda$. Thus, $\mu$ is absolutely continuous with respect to $\lambda$.
    Conversely, if $\mu$ is absolutely continuous with respect to $\lambda$, then by the Lebesgue-Radon-Nikodym theorem, there exists a positive $f\in L^{1}(\mathbb{T},\lambda)$ such that $d\mu=fd\lambda$. So, for any $g\in C(\mathbb{T})$ we have
[…]
    which shows that $\mu$ is a vector state. However, the vector state $\hat{\mu}(g)=\langle g\sqrt{f},~{}\sqrt{f}\rangle_{L^{2}(\lambda)}$ on $L^{\infty}(\mathbb{T},\lambda)$ is an obvious extension of $\mu$ to a weak* continuous functional on $L^{\infty}(\mathbb{T},\lambda)$.
  2. (ii)
    Suppose that $\mu$ has no extension to a PLF $L^{\infty}(\mathbb{T},\lambda)$ that majorize a weak* continuous functional on $L^{\infty}(\mathbb{T},\lambda)$. By part (i), this means that it does not majorize any $\lambda$-absolutely continuous measure on $C(\mathbb{T})$. So, if we consider the Lebesgue decomposition
    $\mu=\mu_{ac}+\mu_{s}$
    of $\mu$ with respect to $\lambda$, we must have $\mu_{ac}=0$, i.e., $\mu=\mu_{s}$. That is $\mu$ is Lebesgue singular with respect to $\lambda$.
    Conversely, let $\mu$ be a Lebesgue singular measure with respect to $\lambda$. Suppose to the contrary that $\mu$ has a PLF extension $\hat{\mu}$ on $L^{\infty}(\mathbb{T},\lambda)$ that majorizes a weak* continuous functional $\gamma$. Consider the Lebesgue decomposition of $\mu$ with respect to $\lambda$ is $\mu=\mu_{ac}+\mu_{s}$. By part (i), $\gamma$ is absolutely continuous, and by our assumption $\mu_{ac}\geq\gamma\neq 0$. Nevertheless, this violates the singularity of $\mu$ with respect to $\lambda$. Thus, no extension of $\mu$ can majorize a non-zero weak* continuous functional on $L^{\infty}(\mathbb{T},\lambda)$.
  3. (iii)
    From part (i), part (ii) and the Lebesgue decomposition theorem.
Therefore, in the classical setting of measure theory and functional analysis, the absolute continuity of a measure is equivalent to the ability of extending the corresponding Riesz-Markov functional to a weak* continuous one. In addition, singularity of measures means the lack of this ability. So, the above theorem justifies the following definition:
Definition 6.
Let $\mu$ and $\lambda$ be two classical, finite, positive, Radon measures on the unit circle $\mathbb{T}$ with their corresponding Riesz-Markov integration functionals on $C(\mathbb{T})$ denoted by $\mu$ and $\lambda$ again.
  1. (i)
    We say that $\mu:C(\mathbb{T})\longrightarrow\mathbb{C}$ is weak* continuous with respect to $\lambda$ if $\mu$ has a weak* continuous extension to a PLF on ${L}^{\infty}(\lambda)$; i.e., there is an extension $\hat{\mu}:{L}^{\infty}({\lambda})\longrightarrow\mathbb{C}$ of $\mu$ which is weak*-continuous and positive. We let $WC[\lambda]$ denote the set of all weak*-continuous measures with respect to $\lambda$.
  2. (ii)
    $\mu$ is called weak*-singular with respect to $\lambda$ if no extension of $\mu$ to ${L}^{\infty}(\lambda)$ can majorize a non-zero weak*-continuous PLF with respect to $\lambda$. We let $WS[\lambda]$ denote the set of all weak*-singular measures with respect to $\lambda$.
Recall that if $\tau$ is a PLF on a unital C*-algebra $A$, then its isotropic left ideal of zero length vectors is $\{a\in A:\tau(a^{*}a)=0\}$.
Definition 7.
Consider two PLFs $\mu$ and $\lambda$ on a unital C*-algebra $A$ with left isotropic ideals $N_{\lambda}\subseteq N_{\mu}$. Now we want to reuse the same classical notation used Remark 4 for the various spaces related to the GNS construction of $\lambda$, e.g., the mapping 4 and Equation 3. Inspired by [20, Theorem III.2.1 and P. 321], we can generalize these notations to denote the GNS representation of $\lambda$ with the cyclic vector $\xi_{\lambda}=1+N_{\lambda}$ by $(\pi_{\lambda},L^{2}(\lambda),\xi_{\lambda})$, the von Neumann algebra generated by the C*-algebra $\pi_{\lambda}(A)$ inside $\mathscr{B}(L^{2}(\lambda))$ by $L^{\infty}(\lambda)=\pi_{\lambda}(A)^{\prime\prime}=\overline{\pi_{\lambda}(A)}^{sot}$, and the predual of $L^{\infty}(\lambda)$ by $L^{1}(\lambda)$. However, note that in the classical setting, $C(\mathbb{T})$ is a closed subspace of $L^{\infty}(\mathbb{T},\lambda)$, and the representation 4 is isometric. Hence, we can extend measures in a well-defined manner without worrying about problems that may arise from changes in the domains. Nevertheless, in the general non-commutative setting, the PLF $\lambda$ is not typically faithful, meaning that $A$ is not isometrically embedded in $L^{\infty}(\lambda)$. Therefore, we need a method to transfer the PLFs $\mu$ and $\lambda$ on $A$ to PLFs on $\pi_{\lambda}(A)$ in a well-defined manner, while ensuring that the main properties of the PLFs are preserved. This is possible by defining:
$\mu^{\prime}(\pi_{\lambda}(a)):=\mu(a)$
and
$\lambda^{\prime}(\pi_{\lambda}(a)):=\lambda(a).$
Recall that in the classical setting, $\pi_{\lambda}$ is an isometric $\ast$-homomorphism, so the above definition covers the classical case as well. Next, we need to show that the above transferred maps are well-defined PLFs.
Proposition 1.
With the notations as in Definition 7, the functionals $\lambda^{\prime}$ and $\mu^{\prime}$ are well defined and give rise to PLFs on $\pi_{\lambda}(A)$. Furthermore, they have positive linear extensions to $L^{\infty}(\lambda)$.
Proof.
By [15, P. 94], the GNS mapping $\pi_{\lambda}:A\longrightarrow\mathscr{B}(L^{2}(\lambda))$ defined by $\pi_{\lambda}(a)(b+N_{\lambda})=N_{\lambda}$ is a $\ast$-homomorphism, and by [15, Theorem 3.1.6] $\pi_{\lambda}(A)$ is a C*-subalgebra of $\mathscr{B}(L^{2}(\lambda))$. Clearly, $\pi_{\lambda}(A)$ embeds isometrically inside $L^{\infty}(\lambda)$. Let $\pi_{\lambda}(a)=0$. So, $\pi_{\lambda}(a)(b+N_{\lambda})=N_{\lambda}$ for any $b\in A$. That is $ab\in N_{\lambda}$. In particular for $b=1$, we have $a\in N_{\lambda}$, so $\lambda(a^{*}a)=0$. By the Kadison inequality [13], we have
$\lambda(a)^{*}\lambda(a)\leq\|\lambda\|\lambda(a^{*}a).$
Thus, $\lambda(a)=0$, so $\lambda^{\prime}$ is well-defined. Since $a\in N_{\lambda}\subseteq N_{\mu}$, we have $\mu(a^{*}a)=0$. By another application of Kadison inequality, we see that $\mu(a)=0$, i.e., $\mu^{\prime}$ is also well-defined. Both $\mu^{\prime}$ and $\lambda^{\prime}$ are PLFs on $\pi_{\lambda}(A)$ since $\pi_{\lambda}$ is a bounded $\ast$-homomorphism between C*-algebras. Note that $\pi_{\lambda}$ is unital since $A$ is unital. Hence, $\|\mu^{\prime}\|=\mu^{\prime}(1)=\mu(1)=\|\mu\|$ by [15, Corollay 3.3.4]. For the second part of the assertion, since $\pi_{\lambda}(A)$ is a C*-sub algebra of $L^{\infty}(\lambda)$, by [15, Theorem 3.3.8] we can extend any PLF on $\pi_{\lambda}(A)$ to a PLF on $L^{\infty}(\lambda)$ with the same norm. ∎
Remark 5.
With Proposition 1 at our disposal, we will drop the prime notations and the homomorphism symbols in expressions like $\mu^{\prime}(\pi_{\lambda}(a))$ and $\lambda^{\prime}(\pi_{\lambda}(a))$, and simply write $\mu(a)$ and $\lambda(a)$, respectively. That is, we regard $\mu(a)$ as $\mu^{\prime}(\pi_{\lambda}(a))$ and vice versa, i.e., $\mu(a)=\mu^{\prime}(\pi_{\lambda}(a))$ and $\lambda(a)=\lambda^{\prime}(\pi_{\lambda}(a))$. However, we know that the actual PLFs $\mu$ and $\lambda$ live on $A$ rather than $\pi_{\lambda}(A)$. Another abuse of notation is that we use the same notation $\mu$ and $\lambda$ for the extended PLFs on $L^{\infty}(\lambda)$ unless otherwise there is a danger of confusion. $\blacksquare$
Now, Remark 4, Theorem 1, Proposition 1, Remark 5, inspire us to generalize Definition 6 from commutative C*-algebras to the non-commutative ones. In addition, we should mention that other main sources for our inspiration were [5, Proposition 2.2], [14, Theorem 10.1.17], [20, Sections III.2 and III.3], [4, Section III.2], and some other results in Kadison and Ringrose’s book, including Proposition 10.1.20, Proposition 10.4.1, and Theorem 10.4.3, in [14].
Definition 8.
Consider two PLFs $\mu$ and $\lambda$ on a unital C*-algebra $A$ such that $N_{\lambda}\subseteq N_{\mu}$. With the notations as in Definition 7 and Remark 5, we say
  1. (i)
    $\mu:A\longrightarrow\mathbb{C}$ is weak* continuous with respect to $\lambda$ if $\mu$ has a weak* continuous extension to a PLF $\hat{\mu}$ on $L^{\infty}({\lambda})$, i.e., the extension $\hat{\mu}:L^{\infty}({\lambda})\longrightarrow\mathbb{C}$ be weak*-continuous. We use $\mu\ll_{w}\lambda$ to show that $\mu$ is weak* continuous with respect to $\lambda$, and we denote the set of all weak*-continuous PLFs with respect to $\lambda$ by $WC[\lambda]$.
  2. (ii)
    Moreover, $\mu$ is called weak*-singular with respect to $\lambda$ if no extension of $\mu$ to $L^{\infty}(\lambda)$ can majorize a non-zero weak*-continuous PLF. We use $\mu\perp_{w}\lambda$ to show that $\mu$ is weak* singular with respect to $\lambda$, and we denote the set of all weak*-singular PLFs with respect to $\lambda$ by $WS[\lambda]$.
  3. (iii)
    The zero functional can be both weak* continuous and weak* singular with respect to $\lambda$.
Remark 6.
The above definition is reasonable since every PLF $\lambda:A\longrightarrow\mathbb{C}$ should be weak* continuous with respect to itself. This is because by the GNS construction we have
$\lambda^{\prime}(\pi_{\lambda}(a))=\lambda(a)=\langle\pi_{\lambda}(a)(1+N_{\lambda}),~{}1+N_{\lambda}\rangle_{L^{2}(\lambda)}$
So, by our notation in Remark 5, we have $\lambda(\pi_{\lambda}(a))=\langle\pi_{\lambda}(a)(1+N_{\lambda}),~{}1+N_{\lambda}\rangle_{L^{2}(\lambda)}$, which is a vector state on $L^{\infty}(\lambda)$ by [15, Theorem 4.2.6]. So, by by [20, Theorem II.2.6] $\lambda$ is indeed a weak* continuous functional on $A$.
Recall that a sub-cone $\mathcal{S}$ of a positive cone $\mathcal{C}$ is hereditary if whenever $s\in\mathcal{S}$ and $c\in\mathcal{C}$ with $c\leq s$, then $c\in\mathcal{S}$.
Theorem 2.
Let $\mu$, $\lambda$ and $\tau$ be three PLFs on a unital C*-algebra $A$ such that $N_{\lambda}\subseteq N_{\mu}$ and $N_{\lambda}\subseteq N_{\tau}$. Then,
  1. (i)
    $\mu$ has a unique weak* Lebesgue decomposition
    $\mu=\mu_{wc}+\mu_{ws}$
    with respect to $\lambda$, where $\mu_{wc}$ and $\mu_{ws}$ are PLFs and $\mu_{wc}$ is the maximal weak* continuous PLF such that $\mu_{wc}\leq\mu$.
  2. (ii)
    $WC[\lambda]$ and $WS[\lambda]$ are hereditary cones.
  3. (iii)
    $(\mu+\tau)_{wc}=\mu_{wc}+\tau_{wc}$ and $(\mu+\tau)_{ws}=\mu_{ws}+\tau_{ws}$
  4. (iv)
    $\mu\leq\tau$ implies $\mu_{wc}\leq\tau_{wc}$ and $\mu_{ws}\leq\tau_{ws}$.
Proof.
  1. (i)
    Note that by a general result in von Neumann algebra theory, see [20, p. 127 and Theorem III.2.14], we have
    $L^{\infty}(\lambda)^{*}=L^{1}(\lambda)\oplus_{1}L^{1}(\lambda)^{\perp},$
    in which, both summands, i.e., the weak* continuous $L^{1}(\lambda)$ and the weak* singular PLFs $L^{1}(\lambda)^{\perp}$, are normed closed sub-spaces of $L^{\infty}(\lambda)^{*}$. In addition, the decomposition into weak* continuous and weak* singular parts is unique. Furthermore, both summands are positive by [14, Theorem 10.1.15-iii].
  2. (ii)
    Note that by [20, p. 127 and Theorem III.2.14] and [14, Theorem 10.1.15-iii], $WC[\lambda]$ and $WS[\lambda]$ are positive cones. So, we need only to prove the hereditary properties. We first show that $WS[\lambda]$ is hereditary. Let $\nu$ be a PLF and $\mu$ be a weak* singular PLF dominating $\nu$. Since $\nu=\nu_{wc}+\nu_{ws}\leq\mu=\mu_{ws}$, we see that $\nu_{wc}\leq\mu$. However, $\mu$ is singular, so by the definition of a singular PLF, we must have $\nu_{wc}=0$. Note that we can prove this in another way. In fact, since $\mu$ is singular, by [20, Theorem III.3.8] we see that for any non-zero projection $e\in L^{\infty}(\lambda)$ there is a non-zero projection $e_{0}\leq e$ such that $\mu(e_{0})=0$. Hence, $\nu_{wc}(e_{0})=0$, which makes $\nu_{wc}$ into a singular PLF. This is only true when $\nu_{wc}=0$. Thus, $\nu=\nu_{ws}$, so $\nu\in WS[\lambda]$.
    Now, let $\nu$ be a PLF and $\mu$ be a weak* continuous PLF dominating $\nu$. Since $\nu=\nu_{wc}+\nu_{ws}\leq\mu=\mu_{wc}$, we see that $0\leq\nu_{ws}\leq\mu_{wc}$. Hence, the PLF $\gamma=\mu_{wc}-\nu_{ws}$ is a PLF. By the uniqueness of the decomposition into weak* continuous and weak* singular parts, see [20, p. 127 and Theorem III.2.14], we obtain that $\gamma_{wc}=\mu_{wc}$ and $\gamma_{ws}=-\nu_{ws}$. However, this means that the singular part of $\gamma$ is negative, which contradicts [14, Theorem 10.1.15-iii].
  3. (iii)
    Note that $\mu+\tau=(\mu+\tau)_{wc}+(\mu+\tau)_{ws}$. Also, $\mu+\tau=(\mu_{wc}+\mu_{ws})+(\tau_{wc}+\tau_{ws})$. So,
    $\mu+\nu=(\mu+\tau)_{wc}+(\mu+\tau)_{ws}=(\mu_{wc}+\tau_{wc})+(\mu_{ws}+\tau_{ws})$
    However, note that the weak* continous and singular PLFs form normed closed sub-spaces of $L^{\infty}(\lambda)^{*}$. Hence, by uniqueness of the decomposition the result follows. Here, the hereditary properties can also be used to show the equality.
  4. (iv)
    For the last assertions note that
    $\mu_{wc}\leq\mu_{wc}+\mu_{ws}\leq\tau_{wc}+\tau_{ws}$
    and
    $\mu_{ws}\leq\mu_{wc}+\mu_{ws}\leq\tau_{wc}+\tau_{ws}.$
    So, $(\tau_{wc}-\mu_{wc})+\tau_{ws}$ and $\tau_{wc}+(\tau_{ws}-\mu_{ws})$ are PLFs, with unique decompositions into positive weak* continuous and positive weak* singular PLFs, see [20, p. 127 and Theorem III.2.14] and [14, Theorem 10.1.15-iii]. Thus, $\tau_{wc}-\mu_{wc}\geq 0$ and $\tau_{ws}-\mu_{ws}\geq 0$.
Now, we want to show that our weak* decomposition method of PLFs is equivalent to other known decomposition theories like Arveson-Gheondea-Kavruk Lebesgue (AGKL) [9]. To do so, we noticed that our splitting PLF $\lambda$ should have some kind of weak tracial property known as the Kubo-Martin-Schwinger (KMS) condition defined below. At the moment, we are unable to relax this condition. We should mention that in the classical setting we are working with abelian C*-algebras, so the KMS condition is automatic. On the other hand, a splitting PLF with the KMS property provides some sharp equations like $(\mu+\nu)_{ac}=\mu_{ac}+\nu_{ac}$, which cannot be achieved within the AGKL theory, see [9, Corollary 3.7-ii]. After proving our equivalence theorem, we can compute the Radon-Nikodym theorem for weak* continuous PLFs.
For a C*-algebra $A$, an automorphism is defined as a $\ast$-homomorphism that is also a bijection, i.e., a $\ast$-isomorphism. The group of all such automorphisms of $A$ is denoted by $\mathrm{Aut}(A)$. In the context of von Neumann algebras, automorphisms are further required to be normal, meaning they preserve the weak*-topology topology. A strongly continuous one parameter group homomorphism is a mapping $\sigma:\mathbb{R}\longrightarrow\mathrm{Aut}(A)$ which is continuous with respect to the Euclidean topology of $\mathbb{R}$ and the strong operator topology topology of $\mathrm{Aut}(A)$. If we put $\sigma_{t}=\sigma(t)$, the continuity of $\sigma$ means that for any $a\in A$ the mapping $\mathbb{R}\longrightarrow A$ given by $t\longrightarrow\sigma_{t}(a)$ is (Euclidean, norm)- continuous. For von Neumann algebras, we require the continuity of the latter mapping when $A$ has the strong operator topology. In this situation, $\sigma$ is called the time evolution of the C*-dynamical (or the W*-dynamical) system $(A,\mathbb{R},\sigma)$, see the Definitions 4 and 5.
The following definition is based on [3, Definition 5.3.1] for the KMS states in quantum mechanics.
Definition 9.
Let $(A,\mathbb{R},\sigma)$ be a C*-dynamical system. A PLF $\lambda$ on a C*-algebra $A$ is called a Kubo-Martin-Schwinger (KMS) PLF for $(A,\mathbb{R},\sigma)$ if
  1. (i)
    for each $a\in A$ the mapping $\mathbb{R}\longrightarrow A$ given by $t\longrightarrow\sigma_{t}(a)$ has an analytic continuation to a mapping $f_{a}:\{z\in\mathbb{C}~{}:~{}0<\mathrm{Im}(z)<1\}\longrightarrow A$ given by $f_{a}(z)=\sigma_{z}(a)$,
  2. (ii)
    the weak tracial property $\lambda(ab)=\lambda(b\sigma_{i}(a))$ holds.
  3. (iii)
    $\lambda$ is time invariant, i.e., $\lambda(\sigma_{t})=\lambda$ for any $t\in\mathbb{R}$
We use the acronym KMS-PLF for a Kubo-Martin-Schwinger positive linear functional.
Remark 7.
Usually one extracts the dynamics $\sigma$ from the PLF $\lambda$. Note that according to [3, Proposition 5.3.3], condition (i) and (ii) imply condition (iii). Also, by [3, Chapter 5] a weaker form of condition (i) is sufficient, i.e., one requires that the members of a (weak*) dense $\ast$-sub-algebra have analytic extensions to the strip $\{z\in\mathbb{C}~{}:~{}0<\mathrm{Im}(z)<\beta\}$, see [3, P. 77] and [16, P. 75]. In this situation, we should replace (ii) by $\lambda(ab)=\lambda(b\sigma_{\beta i}(a))$. The KMS condition puts a restriction on the non commutativity of the PLF $\lambda$ without requiring it to be a full trace. This is important since most C*-algebras do not have a trace, but they can posses a KMS state. Here, $\beta$ measures how far $\lambda$ is being from an actual trace. For $\beta=0$, we have an actual trace. When $\beta\neq 0$, the consensus in physics is to use $\beta=1$, while in mathematics $\beta=-1$ is used because of the coordination with Tomita-Takesaki theory.
Before giving the next example, we need some definitions and constructions related to C*-algebras. The inductive or direct limit of $C^{*}$-algebras is a construction used to build larger $C^{*}$-algebras from a sequence of smaller $C^{*}$-algebras, see [15, Chapter 6].
Definition 10.
Suppose that we have a sequence of C*-algebras $(A_{n})_{n\in\mathbb{N}}$, and assume that when $n\leq m$ there is $\ast$-homomorphism $\phi_{n,m}:A_{m}\to A_{n}$. Define
$A^{\prime}=\{(a_{n})\in\prod_{n\in\mathbb{N}}A_{n}~{}:~{}\exists~{}K\in\mathbb{N}~{}\forall n\geq m\geq K,~{}a_{n}=\phi_{n,m}(a_{m})\}$
Since $\ast$-homomorphisms between C*-algebras are norm decreasing, for any $a=(a_{n})_{n\in\mathbb{N}}\in A^{\prime}$ the sequence of norms $(\|a_{n}\|)_{n\in\mathbb{N}}$ is eventually decreasing. Hence, the limit
$p(a)=\lim_{n\to\infty}\|a_{n}\|$
[…]
The following example is taken from [2, P. 13] and [16, P. 77], which is important in quantum spin systems.
Example 1.
Consider the lattice $\mathbb{Z}^{n}$, and suppose that to each point of $x\in\mathbb{Z}^{n}$ a copy of a finite dimensional Hilbert space $\mathcal{H}_{x}$ is attached. For a finite family $\mathcal{F}$ in $\mathbb{Z}^{n}$, we form the tensor product $\mathcal{H}_{\mathcal{F}}=\otimes_{x}\mathcal{H}_{x}$ and the operator algebra $A(\mathcal{F})=\mathscr{B}(\mathcal{H}_{\mathcal{F}})$. The direct limit of $A(\mathcal{F})$, denoted by $\mathcal{A}$, is called the C*-algebra of local observable, [16, P. 56]. Note that $\mathcal{A}$ is a sub-algebra of a suitable $\mathscr{B}(\mathcal{H})$, where $\mathcal{H}$ is related to the infinite tensor product of Hilbert spaces. One can show that there a closed unbounded operator $H$, known as the Hamiltonian of the system, such that $e^{-\beta H}$ is a trace class operator and $\lambda(a)=\mathrm{tr}(ae^{-\beta H})$ is a KMS-PLF on $\mathcal{A}$. Note that the domain $\mathcal{D}$ of $H$ is dense in the Hilbert space $\mathcal{H}$, and $e^{-\beta H}$ is just invertible on this dense domain not all of $\mathcal{H}$. Thus, being a trace class for $e^{-\beta H}$ does not violate the fact that the ideal of trace class operators contains no invertible everywhere defined operator, see [2, P.85]. Such an operator is called a density matrix for $\lambda$, see [2, P. 76]. Also, $\sigma_{t}(a)=e^{itH}ae^{-itH}$ determines the dynamics of the system, see [3, P.76].
Example 2.
Let $\mathscr{M}$ be a von Neumann algebra inside some $\mathscr{B}(\mathscr{H})$, and suppose that $\mathscr{M}$ has a separating and cyclic unit vector $\eta$. Then, by Tomita-Takesaki theory [21], one can show that the vector state $\phi(a)=\langle a\eta,~{}\eta\rangle$ is a KMS state, see [2, P. 285]. In this situation, we let $S$ be the closure of the densely defined operator
$S_{0}:\mathscr{M}\eta$$\longrightarrow \displaystyle\mathscr{H}$
$m\eta$$\longrightarrow$ $m^{*}\eta,$
and we put $\Delta=S^{*}S$. By the Tomita-Takesaki theory [21], the time evolution is given by (see [2, P. 284])
$\sigma:\mathbb{R}$$\longrightarrow$ $\mathrm{Aut}(\mathscr{M})$
$\sigma_{t}(a)$ $=$ $\Delta^{it}a\Delta^{-it}~{}.\blacksquare$
Let’s recall some other forms of absolute continuity and singularity for PLFs. In the classical measure theory, absolute continuity of measures on $\mathbb{T}$ can be translated into the language of convergence of measures. Namely, given two finite positive regular Radon measures $\mu$ and $\lambda$ on $\mathbb{T}$, one can show that $\mu$ is absolutely continuous (AC) with respect to $\lambda$, in notations $\mu\ll\lambda$, if there is an increasing sequence of finite positive regular Radon measures $\mu_{n}$, which are dominated by $\lambda$, and increasing monotonically to $\mu$:
$0\leq\mu_{n}\leq\mu,\quad\mu_{n}\uparrow\mu,$
$\forall~{}n~{}\exists~{}t_{n}>0,\quad\mu_{n}\leq t_{n}\lambda.$
The following definition, provided in [9], generalizes the above concept to the non-commutative setting.
Definition 11.
Consider two PLFs $\mu$ and $\lambda$ on a unital C*-algebra $A$ such that $N_{\lambda}\subseteq N_{\mu}$.
  1. (i)
    We say $\mu$ is absolutely continuous with respect to $\lambda$ in the sense of Gheondea-Kavruk (GK), denoted by $\mu\ll\lambda$, if
    • there is an increasing sequence $(\mu_{n})_{n\in\mathbb{N}}$ of PLFs such that $\mu_{n}\uparrow\mu$ point-wise, i.e., in the weak* topology of $A^{*}$
    • for any $n\in\mathbb{N}$, there is a positive $t_{n}$ such that $\mu_{n}\leq t_{n}\lambda$
    We denote the set of all absolutely continuous PLFs with respect to $\lambda$ by $AC[\lambda]$.
  2. (ii)
    Moreover, $\mu$ is called singular with respect to $\lambda$, denoted by $\mu\perp\lambda$, if the only PLF majorized by $\mu$ and $\lambda$ is the zero PLF. We denote the set of all singular PLFs with respect to $\lambda$ by $SG[\lambda]$.
Theorem 3.
Let $(A,\mathbb{R},\sigma)$ be a C*-dynamical system. Assume that $\mu$ and $\lambda$ are two PLFs on it with $N_{\lambda}\subseteq N_{\mu}$. Suppose further that $\lambda$ is a KMS-PLF. Then,
  1. (i)
    $\mu\ll_{w}\lambda$ if and only if $\mu\ll\lambda$, and
  2. (ii)
    $\mu\perp_{w}\lambda$ if and only if $\mu\perp\lambda$.
Proof.
  1. (i)
    Suppose that $\mu\ll_{w}\lambda$. By Definition 8, $\mu$ has a weak* continuous extension to a PLF on $L^{\infty}({\lambda})$. By [20, P. 156], there is a sequence $(\xi_{n})_{n\in\mathbb{N}}$ of vectors in $L^{2}(\lambda)$ such that $s=\sum_{n=1}^{\infty}\|\xi_{n}\|^{2}<\infty$ and $\mu(y)=\sum_{n=1}^{\infty}\langle y\xi_{n},~{}\xi_{n}\rangle_{\lambda}$. First of all, we claim that we can choose $\xi_{n}\in\frac{A}{N_{\lambda}}$. For any $n$, there is a sequence $(\xi_{n,k})_{k}$ in $A$ such that $\xi_{n}=\lim\limits_{k\rightarrow\infty}\xi_{n,k}+N_{\lambda}$ in $L^{2}(\lambda)$. In particular, $\|\xi_{n}\|^{2}=\lim\limits_{k\rightarrow\infty}\lambda(\xi_{n,k}^{*}\xi_{n,k})$. Thus, for any $n$ and any $\epsilon>0$, there is a $K=K(n,\epsilon)$ such that if $n\geq K$, then
    $\|\xi_{n}\|^{2}-\frac{\epsilon}{2^{n}}<\lambda(\xi_{n,k}^{*}\xi_{n,k})<\|\xi_{n}\|^{2}+\frac{\epsilon}{2^{n}}$ (4)
    For any natural $N$, we use induction on to extract a suitable sequence $(\zeta_{N})$ in $\frac{A}{N_{\lambda}}$ from $(\xi_{n,k})_{k}$. By the above equation, within the $\frac{\epsilon}{2^{k}}$-neighborhood of $\xi_{k}$, we can find infinitely many element of $\frac{A}{N_{\lambda}}$ close to $\xi_{n}$. For $N=1$, we can choose an element $\zeta_{1}$ within $\frac{\epsilon}{2}$-neighborhood of $\xi_{1}$. For $N=2$, we can choose an element $\zeta_{2}$ within $\frac{\epsilon}{2^{2}}$-neighborhood of $\xi_{2}$, and different from $\zeta_{1}$. Clearly for any finite $N$, we can proceed by induction and choose an element $\zeta_{N}$ within $\frac{\epsilon}{2^{N}}$-neighborhood of $\xi_{N}$, and different from the previous ones. The next step, $N+1$, should be obvious by now. By equation 4, we have
    $\|\xi_{N}\|^{2}-\frac{\epsilon}{2^{N}}<\lambda(\zeta_{N}^{*}\zeta_{N})=\|\zeta_{N}\|^{2}<\|\xi_{N}\|^{2}+\frac{\epsilon}{2^{N}}$ (5)
    So, $\sum_{N=1}^{\infty}\|\zeta_{N}\|^{2}=s<\infty$. Hence, without loss of generality, we can assume that in $\mu(y)=\sum_{k=1}^{\infty}\langle y\xi_{k},~{}\xi_{k}\rangle_{\lambda}$, we have $\xi_{k}=x_{k}+N_{\lambda}\in\frac{A}{N_{\lambda}}$. However,
    $\langle y\xi_{k},~{}\xi_{k}\rangle_{\lambda}=\langle yx_{k}+N_{\lambda},~{}x_{k}+N_{\lambda}\rangle_{\lambda}=\lambda(x_{k}^{*}yx_{k});$
    hence, $\mu(y)=\sum_{k=1}^{\infty}\lambda(x_{k}^{*}yx_{k})$. Define a sequence of linear functionals by
    $\mu_{n}(y)=\sum_{k=1}^{n}\lambda(x_{k}^{*}yx_{k}),\quad\forall~{}n\in\mathbb{N}$
    We show that $\mu_{n}$’s are the desired PLFs, which satisfy the conditions of absolute continuity in Definition 2. When $y$ is positive, i.e., $y=a^{*}a$ for an element $a\in A$, we see that $\mu_{n}(y)=\sum_{k=1}^{n}\lambda((ax_{k})^{*}(ax_{k}))$ is positive. Furthermore, it is obvious from the construction that $\mu_{n}\uparrow\mu$. So, for any $n\in\mathbb{N}$, we only need to find a $t_{n}>0$ such that $\mu_{n}\leq t_{n}\lambda$ on positive elements like $y=a^{*}a$.
    • Suppose that $\lambda(y)=0$. So, $a\in N_{\lambda}\subseteq N_{\mu}$; thus, $a\in N_{\mu}$ and $\mu(y)=\mu(a^{*}a)=0$. In particular, $\mu_{n}(y)=0$. So, in this case, the inequality $\mu_{n}\leq t_{n}\lambda$ holds for any positive $t_{n}$.
    • For the case $\lambda(y)>0$, define a bi-linear form $V$ on $A$ by
      $V:A\times A$$\longrightarrow \displaystyle\mathbb{C}$
      $V(a,b)$ $=$ $\frac{\lambda(bya)}{\lambda(y)}$
      Recall that $\lambda$ is a KMS-PLF, so for $u=b$ and $v=ya$, we have $\lambda(uv)=\lambda(v\sigma_{i}(u))$; hence,
      $V(a,b)=\frac{\lambda(bya)}{\lambda(y)}=\frac{\lambda(ya\sigma_{i}(b))}{\lambda(y)}=\frac{\lambda\cdot y(a\sigma_{i}(b))}{\lambda\cdot y(1)}$
      Since $y$ is positive, the functional $\lambda\cdot y$ is positive, and by Remark 3,
      $\|\lambda\cdot y\|=\lambda\cdot y(1)=\lambda(y).$
      Therefore,
      $|V(a,b)|\leq\frac{\|\lambda\cdot y\|\|a\sigma_{i}(b))\|}{|\lambda(y)|}=\|a\sigma_{i}(b))\|\leq\|a\|\|b\|$
      Consequently, $V$ is a bounded bi-linear form on the C*-algebra $A$. By Generalized Grothendieck inequality [10], we can find states $\phi_{i}$,  $i=1,\cdots,4$ , on $A$ such that
      $|V(a,b)|\leq[\phi_{1}(a^{*}a)+\phi_{2}(aa^{*})]^{\frac{1}{2}}[\phi_{3}(b^{*}b)+\phi_{4}(bb^{*})]^{\frac{1}{2}}$
      Applying this to $V(x_{k},x_{k}^{*})=\frac{\lambda(x_{k}^{*}yx_{k})}{\lambda(y)}$, we get
      $|\frac{\lambda(x_{k}^{*}yx_{k})}{\lambda(y)}|\leq[\phi_{1}(x_{k}^{*}x_{k})+\phi_{2}(x_{k}x_{k}^{*})]^{\frac{1}{2}}[\phi_{3}(x_{k}x_{k}^{*})+\phi_{4}(x_{k}^{*}x_{k})]^{\frac{1}{2}}\leq 2\|x_{k}\|^{2}$
      Thus, $|\lambda(x_{k}^{*}yx_{k})|\leq 2\|x_{k}\|^{2}\lambda(y)$. So,
      $\mu_{n}(y)=\sum_{k=1}^{n}\lambda(x_{k}^{*}yx_{k})\leq\left(2\sum_{k=1}^{n}\|x_{k}\|^{2}\right)\lambda(y).$
    So, in both cases above, by putting $t_{n}=\left(2\sum_{k=1}^{n}\|x_{k}\|^{2}\right)$, we see that $\mu_{n}\leq t_{n}\lambda$. Hence, second condition of the definition of absolute continuity is satisfied. Thus, $\mu\ll\lambda$.
    On the other hand, assume that $\mu\ll\lambda$. By the Definition 11, there is an increasing sequence $(\mu_{n})_{n\in\mathbb{N}}$ of PLFs such that $\mu_{n}\uparrow\mu$, and for any $n\in\mathbb{N}$ there is a positive $t_{n}$ such that $\mu_{n}\leq t_{n}\lambda$. Since the von Neumann algebra $L^{\infty}(\lambda)$ is generated by its positive elements, we only need to check the weak* continuity of any functional on $L^{\infty}(\lambda)_{+}$. So, assume that $(y_{\alpha})$ is a net of positive elements converging to zero in the weak* topology. By the GNS construction, $\lambda$ is weak* continuous with respect to itself, and by $\mu_{n}\leq t_{n}\lambda$, we deduce that $\mathrm{lim}_{\alpha}\mu_{n}(y_{\alpha})=0$. That is, each $\mu_{n}$ is a normal PLF. In addition, the sequence of normal PLFs $(\mu_{n})$ converges to the PLF $\mu$ in the weak* topology, so by [20, Theorem III.5.1], the PLF $\mu$ must be normal; hence, $\mu\ll_{w}\lambda$.
  2. (ii)
    Let $\mu\perp_{w}\lambda$, but $\mu\not\perp\lambda$. By [9], this means that we can decompose $\mu$ against $\lambda$, i.e., $\mu=\mu_{ac}+\mu_{s}$, where $\mu_{ac}\neq 0$ and $\mu_{ac}\ll\lambda$. By part (i), we conclude that $\mu_{ac}$ is weak* continuous with respect to $\lambda$. Since $\mu_{ac}\leq\mu$, we see that $\mu$ dominates the non-zero weak* continuous PLF $\mu_{ac}$. However, this is impossible as $\mu$ is weak* singular.
    The reverse implication is proved similarly. Indeed, if $\mu\perp\lambda$ but $\mu\not\perp_{w}\lambda$, then we have a weak* decomposition $\mu=\mu_{wc}+\mu_{ws}$, where $\mu_{wc}\neq 0$. Now, $\mu_{wc}\leq\mu$ and $\mu_{wc}\ll\lambda$ by part (i). So, from Definition 11 we can find a non-zero PLF $\rho$ such that $\rho\leq\mu$ and $\rho\leq\lambda$. However, this contradicts the fact that $\mu\perp\lambda$.
We are now in a position to strengthen some of the results of [9] for splitting KMS-PLFs.
Theorem 4.
Let $(A,\mathbb{R},\sigma)$ be a C*-dynamical system. Assume that $\mu$, $\nu$ and $\lambda$ are PLFs on it with $N_{\lambda}\subseteq N_{\mu}$ and $N_{\lambda}\subseteq N_{\nu}$. Suppose further that $\lambda$ is a KMS-PLF. Then,
  1. (i)
    $AC[\lambda]$ is a positive hereditary cone,
  2. (ii)
    $SG[\lambda]$ is a positive hereditary cone,
  3. (iii)
    $(\mu+\nu)_{ac}=\mu_{ac}+\nu_{ac}$ with respect to $\lambda$,
  4. (iv)
    $(\mu+\nu)_{s}=\mu_{s}+\nu_{s}$ with respect to $\lambda$,
  5. (v)
    If $\mu\leq\nu$, then $\mu_{ac}\leq\nu_{ac}$ and $\mu_{s}\leq\nu_{s}$
  6. (vi)
    In particular, the weak* Lebesgue decomposition $\mu=\mu_{wc}+\mu_{ws}$ with respect to $\lambda$ gives rise to the maximal Arveson-Gheondea-Kavruk Lebesgue (AGKL) decomposition $\mu=\mu_{ac}+\mu_{s}$ with respect to $\lambda$. Conversely, the maximal AGKL decomposition implies the weak* Lebesgue decomposition.
Proof.
The proof is a result of Theorem 2 and the equivalent Theorem 3. ∎
Finally, we can use the result of [9] to find the Radon-Nikodym derivative for the weak* continuous part of the weak*-Lebesgue decomposition of $\mu$ with respect to $\lambda$. Let $D$ be an unbounded operator on a Hilbert space $\mathscr{H}$, and $\mathscr{M}$ be a von Neumann algebra inside $\mathscr{B}(\mathscr{H})$. We say that $D$ is affiliated with $\mathscr{M}$, and we write $D\sim\mathscr{M}$ if for any $B\in\mathscr{M}^{\prime}$, we have $BD\subseteq DB$; i.e., $\mathrm{Dom}(D)$ is $B$-invariant, and $BDh=DBh$ for any $h\in\mathrm{Dom}(D)$. What is important here is that $D$ commutes with elements of $\mathscr{M}^{\prime}$ over the domain of $D$. When, $D$ is bounded, the condition $D\sim\mathscr{M}$ means $D\in\mathscr{M}$.
Theorem 5.
Let $(A,\mathbb{R},\sigma)$ be a C*-dynamical system. Assume that $\mu$ and $\lambda$ are PLFs on it with $N_{\lambda}\subseteq N_{\mu}$, and $\lambda$ is a KMS-PLF. If $\mu\ll_{wc}\lambda$, then there is a unique and possibly unbounded operator $D_{\lambda}\mu$ such that
  1. (i)
    $D_{\lambda}\mu$ is positive.
  2. (ii)
    $D_{\lambda}\mu$ is affiliated with the von Neumann algebra $L^{\infty}(\lambda)^{\prime}$.
  3. (iii)
    $\pi_{\lambda}(A)\xi_{\lambda}$ is a subset of the domain of $\sqrt{D_{\lambda}\mu}$ Here, $\xi_{\lambda}=1+N_{\lambda}$ is the cyclic vector of the GNS representation of $\lambda$.
  4. (iv)
    $\mu(a)=\langle\pi_{\lambda}(a)\sqrt{D_{\lambda}\mu}~{}\xi_{\lambda}~{},~{}\sqrt{D_{\lambda}\mu}~{}\xi_{\lambda}\rangle_{\lambda}$.
  5. (v)
    If $\mu\leq t\lambda$ for a positive number $t\in\mathbb{R}$, then $D_{\lambda}\mu$ is a bounded positive operator which belongs to $L^{\infty}(\lambda)^{\prime}$ and $\mu(a)=\langle\pi_{\lambda}(a)D_{\lambda}(\mu)(\xi_{\lambda})~{},~{}\xi_{\lambda}\rangle_{\lambda}$. Also, the assignment $\mu\longrightarrow D_{\lambda}(\mu)$ is an affine mapping with respect to $\mu$.
Proof.
The proofs of $i$ to $iv$ is a result of the equivalent Theorem 3 and [9, 2.11]. The final part is a result of the equivalent Theorem 3 and [9, Corollary 2.3]. ∎
Remark 8.
When $\mu\ll_{wc}\lambda$, by Theorem 3 and Definition 11, we know that there is an increasing sequence $(\mu_{n})_{n\in\mathbb{N}}$ of PLFs such that $\mu_{n}\uparrow\mu$ point-wise, i.e., in the weak* topology of $A^{*}$. Also, for any $n\in\mathbb{N}$ there is a positive $t_{n}$ such that $\mu_{n}\leq t_{n}\lambda$. The latter condition implies that $\mu_{n}\ll_{ac}\lambda$, and hence each $D_{\lambda}\mu_{n}$ is a positive bounded operator in $L^{\infty}(\lambda)^{\prime}$. By [9, Part F], one finds that $(I+D_{\lambda}\mu_{n})^{-1}\xrightarrow{SOT}(I+D_{\lambda}\mu)^{-1}$. This convergence is described by saying that $D_{\lambda}\mu_{n}\xrightarrow{SRS}D_{\lambda}\mu$, where “SRS” stand for the strong resolvent sense. Note that the number $-1$ does not belong to the spectra of $D_{\lambda}\mu_{n}$’s, so by [18, Theorem 5.1.9] the operators $I+D_{\lambda}\mu_{n}$ and $I+D_{\lambda}\mu$ are invertible. In summary, though the Radon-Nikodym derivative $D_{\lambda}\mu$ is generally an unbounded operator, one can still approximate it in a suitable sense by bounded Radon-Nikodym derivatives.
For the following remark, we do not define the new terms and refer the reader to the cited references for further details.
Remark 9.
Let $\mu$ and $\lambda$ be positive linear functionals defined on the Cuntz-Toeplitz operator system, as in [11] and [17]. One can apply the method presented here to decompose $\mu$ with respect to $\lambda$. First, as in [17], $\lambda$ can be classified into two cases: Cuntz and non-Cuntz. Second, it is essential to ensure that the KMS condition for $\lambda$ holds on the ambient Cuntz C*-algebra. In the non-Cuntz case, the decomposition method of [17] generalizes the approach from [11]. In the Cuntz case, $\lambda$ has a unique extension to the ambient Cuntz C*-algebra by [12, Proposition 5]. This allows the decomposition to be carried out in that setting, yielding the desired result. While there are technical subtleties that must be addressed, the key missing component in [11] and [17] is the KMS condition. In fact, by Example 2 and [7, Example 1.5.8], the Lebesgue expectation $m$ naturally satisfies the KMS condition, and it is this hidden property that makes the decomposition and the equivalence in [11] possible.